1. The enthalpy change for the following reaction is 81.1 kJ. Using bond energies, estimate the H-Cl bond energy in HCl(g).

2HCl(g) + Br2(g)2HBr(g) + Cl2(g)

___ kJ/mol

2. The enthalpy change for the following reaction is -137 kJ. Using bond energies, estimate the C-C bond energy in C2H6(g).

C2H4(g) + H2(g)C2H6(g)

___ kJ/mol
1
Expert's answer
2019-11-26T06:39:34-0500
  1. "2HCl(g) + Br_2(g) \rightarrow 2HBr(g) + Cl_2(g)" "\Delta H_{rxn} = 81.1 kJ"
"\Delta H _{rxn} = \sum \Delta H(bonds broken) - \sum \Delta H (bonds formed)"

Bonds energies:

Br-Br 193 kJ/mol

H-Br 366 kJ/mol

Cl-Cl 242 kJ/mol



"81.2 = [2\times(H-Cl)+1\times (Br-Br)] - [2\times(H-Br)+1\times(Cl-Cl)]"


"81.2 = [2\times(H-Cl)+193]-[2\times366+242]"

Bond energy of H-Cl is 431.1 kJ/mol


2.

"C_2H_4(g) + H_2(g) \rightarrow C_2H_6(g)" "\Delta H_{rxn} = -137 kJ"






"\Delta H_{rxn} = \sum \Delta H (bonds broken) - \sum \Delta H(bonds formed)"

Bonds energies:

C=C 614 kJ/mol

C-H 413 kJ/mol

H-H 436 kJ/mol


"-137 = [1\times (C=C)+ 4\times(C-H)+1\times(H-H)]-[1\times(C-C)+6\times(C-H)]"


"-137 = [614+4\times413+436]-[(C-C)+6\times413]"


Bond enrgy of C-C is: 361 kJ/mol